Parker Plastic, Inc., manufactures plastic mats to use with rolling office chairs. Its standard cost information for last year follows:

Standard Quantity Standard Price (Rate) Standard Unit Cost
Direct materials (plastic) 11 sq ft. $ 1.26 per sq. ft. $ 13.86
Direct labor 0.3 hr. $ 11.70 per hr. 3.51
Variable manufacturing overhead (based on direct labor hours) 0.3 hr. $ 2.10 per hr. 0.63
Fixed manufacturing overhead $441,600 ÷ 920,000 units) 0.48


Parker Plastic had the following actual results for the past year:

Number of units produced and sold 1,140,000
Number of square feet of plastic used 11,900,000
Cost of plastic purchased and used $ 15,470,000
Number of labor hours worked 318,000
Direct labor cost $ 3,688,800
Variable overhead cost $ 1,030,000
Fixed overhead cost $ 375,000

Calculate Parker Plastic’s direct materials price and quantity variances. (Do not round intermediate calculations. Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable.)

Direct Material Price Variance
Direct Material Quantity Variance

with F or UF
Answer #1

Materials Price Variance = (SP-AP) * AQ

= (1.26 - 15,470,000/11,900,000) * 11,900,000

= 476,000 U

Materials Quantity Variance = (SQ - AQ) * SR

= (11*1,140,000 - 11,900,000) * 1.26

= 806,400 F

COMMENT IF YOU FACE ANY ISSUES WILL GET BACK TO YOU IMMEDIATELY